# The Value of Information in Human-AI Decision-making

### v1

@@ 5.2 Results
How much decision-relevant information do each agent’s decisions offer? We first compare the information value of the AI predictions to the decision problem to that of the human decisions in the first round (without AI assistance). To contextualize the value of each, we first construct a scale ranging from the expected payoff of rational DM with no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, to that of the rational agent who has access to all information, i.e., $\mathrm{R}^{\pi,S}(\Sigma_{1}\cup\ldots\cup\Sigma_{n}\cup D^{\text{Human}}\cup D^{\text{Human-AI}}\cup D^{\text{AI}})$. The lower-bound represented by the rational DM with no information is $0.75$ in Brier score, which is equivalent to the payoff achieved by a random guess drawn from $\text{Bernoulli}(0.5)$.
@@
Using this scale, Figure 2(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no information baseline, while human decisions only provide about $15\%$. Hence, human decisions are only weakly informative for the problem.
@@
We next consider the human-AI decisions. Given that the AI predictions contained a significant portion of the total possible information value, we might hope that when participants are provided with AI predictions, their performance comes close to the full information baseline. However, the information value contained in human-AI decisions is also only take a small proportion of the total possible information value ($30\%$). This aligns with the findings by Guo et al. (2024) that humans tend to rely on the AI but are bad at distinguishing when AI predictions are correct.
@@
How much additional decision-relevant information do the available features offer over each agent’s decisions? To understand what information might improve human decisions, we assess the $ACIV$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the exisiting information in the agents’ decisions. As shown on the fifth row in Figure 2, the presence of a flickering face offers larger $ACIV$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 2, the presence of a individual with dark skin offers larger $ACIV$ over AI predictions than over human decisions. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
@@
By comparing the $ACIV$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to human did not help human-AI to exploit the observed signals in their decisions. As shown in Figure 2, with the assistance of AI, the human-AI teams’ decisions have similar $ACIV$ compared to human decisions (note that the $ACIV$ over human decisions shows significant difference on same signal from the $ACIV$ over AI predictions), except a little improvement on the presence of a flickering face. This finding further comfirms the hypothsis on human simply relying on AI predictions without processing the information contained in them.
@@ 6 Experiment III: Information-based Local Explanation on Recidivism Prediction
We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendent. We apply the framework to evaluate and augment saliency-based explanations.
Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation. (C) The waterfall plot of an example of SHAP on instance 4 and (D) the waterfall plot of an example of ILIV-SHAP on instance 4. See more examples of ILIV-SHAP in Appendix C.
@@ 6.1 Data and Model
We use the dataset from COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendents with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. (2020) to represent the human decisions, which contains recidivism decisions on a subset of instances from Angwin et al. (2022)’s experiment on laypeople [^4] Here the experimental decisions avoid the problem of performative prediction (Perdomo et al., 2020). Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model (Chen and Guestrin, 2016) on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. (2020).
We use Brier score as the payoff function, with the binary payoff-related state $\omega$ indicating whether the defendent gets rearrested within two years. We use the features of defendent contained in COMPAS dedaset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrest and charges for several offense categories) and the predicted score from COMPAS system [^5] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, signals, human decisions and AI predictions.
Constructing instance-level signals.
Denote the XGBoost model’s predictions as a random variable $Y^{pred}=f(X)$, where $X$ denotes the random variable of data features and $f(\cdot)$ denotes the model’s predictive function. We construct $Z_{f(x)}$ as $\mathbb{1}[Y^{pred}=f(x)]$ for every $x$ and use it to calculate the instance-level agent-complementary information value ($ILIV$) of $Y^{pred}$ on instances $x$ (Definition 3.3). We generate the information-based explanations by ILIV-SHAP in this demonstration.
@@ 6.2 Results
How well does SHAP explain complementary information offered by the AI prediction over what humans already know? Figure 3 (A) and (B) compares the distribution of feature-attribution scores from SHAP to those from our ILIV-SHAP. We observe multiple discrepancies in feature importance across the methods. For instance, in Figure 3(A), the age feature negatively correlates with the AI’s prediction, indicating that younger defendants tend to receive higher predictions. However, Figure 3(B) ILIV-SHAP shows no such association between age and the AI’s information value over human decisions. This implies that, although age influences the AI’s prediction, on average it does not provide additional information that humans lack.
Conversely, some features that correlate strongly with the $ILIV$ of the AI prediction are relatively unimportant for the AI’s raw prediction. For example, in Figure 3(A), low decile scores predicted by the COMPAS (colored in blue) have a relatively small impact on the AI’s numerical output compared with other features. In contrast, Figure 3(B) reveals that some low decile scores yield large positive $ILIV$, suggesting that this feature might offer valuable information for human decision-makers despite not significantly changing the AI’s raw prediction.
How can an information-based explanation enhance understanding of AI predictions? We argue that an information-based explanation (via ILIV-SHAP) can serve as a supplement to saliency-based explanations, providing users a sense of whether the individual prediction can help them make decisions. First, an information-based explanation conveys whether the prediction correlates with the payoff-related state. For example, Figure 3(C) shows that the model prediction $f(x)$ deviates $+0.383$ from the prior prediction $\text{\bf E}\![{f(X)}]$ on instance 4. However, the information-based explanation of the same instance in Figure 3(D) shows that the $ILIV$ of the AI prediction over human decisions does not change much from the $ILIV$ of the prior prediction (which is very low at $0.005$). This suggests to the user that focusing on the AI prediction on this instance is not necessary to good decision-making.
Second, the information-based explanation conveys to the user whether the AI gets useful information beyond their own information from a certain feature. In Figure 3(A), the defendent’s prior count record (priors_count=3) changes the AI prediction the most ($+0.186$). However, Figure 3(D) shows that priors_count = 3 contributes marginally to the $ILIV$ ($+0.00004$). Therefore, ILIV-SHAP conveys that even if priors_count=3 changes the AI prediction significantly, the AI does not necessarily better predict the payoff-related state as a result of it.
@@ 7 Discussion and Limitations
We propose a decision-theoretic framework for assigning value to information in human-AI decision-making. Our methods quantify the additional information value of any signals over an agent’s decisions. The three demonstrations show how quantified information value can be used in model selection, empricial evaluation, and explanation design. These demonstrations are just a few of many possible uses cases. For example, alternative explanation approaches could be compared via their information value to human decisions. Information value analysis could drive elicitation of human private signals or decision rules to further improve the pairing. New explanation strategies could integrate visualized information value with conventional depictions of feature importance.
Information value cannot definitively establish that particular signals were used by a human. It is always possible that the human has other private signals offering equivalent information to a feature being analyzed. Our framework cannot account for private unobservable signals even though they might have strong correlation with the payoff-related state and agent deicions. However, quantifying the value of observed information is a tool toward learning about information that may exist beyond a problem definition.
Our framework quantifies the best-attainable performance improvement from integrating signals in decisions. This does not necessarily mean that empirically those signals will lead to agents better performing. However, the motivating idea is that Bayesian decision theory provides a theoretical basis that can be adapted to support informative comparisons to human behavior. For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the Bayesian decision-maker to only make monotonic decisions with AI prediction and their own predictions.
@@ Appendix A The Combinatorial Nature of the Value of Signals
$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(ACIV(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (1)
@@
The following algorithm provides a polynomial-time approximation of the Shapley value of $ACIV$. Under the assumption of submodularity, it orders the signals the same as the Shapley value.
Algorithm 1 Algorithm 1
@@
4:     $\phi^{\prime}_{j}=ACIV(\Sigma_{j};V^{*})\ for\ each\ j$
@@
Appendix B Robustness analysis in Experiment I
Figure 4: Robust analysis for experiment I on all V-shaped payoff functions. The kink $\mu$ is shown on the x-axis. Each subplot displays the difference between the $ACIV$ on the row model over human decisions and the $ACIV$ on the column model over human decisions. The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
Appendix C More examples of ILIV-SHAP

### v3

@@ 5.2 Results
How much decision-relevant information do agents’ decisions offer?
@@
We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance). Figure 2(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
@@
We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$). This is consistent with the findings of Guo et al. (2024) that humans are bad at distinguishing when AI predictions are correct.
How much additional decision-relevant information do the available features offer over agents’ decisions?
@@
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions. As shown on the fifth row in Figure 2, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 2, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
@@
By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions. As shown in Figure 2, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face. This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
@@ 6 Experiment III: Information-based Local Explanation on Recidivism Prediction
We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set. (C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18. The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output). See further comparisons between SHAP and ILIV-SHAP in Appendix D.
@@ 6.1 Data and Model
We use the COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. (2020) to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al. (2022)’s experiment on laypeople. Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model (Chen and Guestrin, 2016) on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. (2020).
We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism (Green and Chen, 2019a; Lin et al., 2020; Fogliato et al., 2021). The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years. We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^4] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
@@ 6.2 Results
SHAP explanations do not communicate the complementary information of AI predictions.
Specifically, the changes in information value of AI predictions caused by marginalizing out features, i.e., ILIV-SHAP scores, are not predictable from the changes in AI predictions, i.e., SHAP scores, and the feature values. Figure 3 (A) and (B) compare the distribution of feature importance scores from SHAP to those from ILIV-SHAP. The scores of prior_counts (the number of prior criminal charges), age, and decile_score (the score from the COMPAS system), which are three features with the largest scores in both SHAP and ILIV-SHAP, show different correlations with the feature values in SHAP and ILIV-SHAP. For instance, in Figure 3(A), most of the low values (blue dots) of prior_counts have a negative SHAP score, and most of the high values (red dots) have a positive SHAP score. This indicates that when the value of prior_counts is low, marginalizing out prior_counts from the features leads to a higher expected model prediction than not marginalizing out, and vice versa. However, this correlation does not hold for ILIV-SHAP. In Figure 3(B), the high values of prior_counts appear on both the left side (below zero) and the right side (above zero), indicating having low ILIV-SHAP scores and high ILIV-SHAP scores at the same time. This mismatch between the distributions of SHAP scores and ILIV-SHAP scores implies that the impact of features on the information value cannot be directly inferred (at least monotonically) from their impact on model output and their values.
ILIV-SHAP can augment SHAP on communicating the complementary information of AI predictions.
We argue that ILIV-SHAP can supplement saliency-based explanations to provide users with a sense of how the AI predictions can help improve their decisions. With ILIV-SHAP, users can check whether the prediction contains decision-relevant information by looking at the total $\operatorname{ILIV}$ of the AI prediction, which indicates the maximum improvement in expected payoff that could be achieved by incorporating this prediction. Figure 3(D) and (F) show two different cases, where the AI predictions contain substantial human-complementary information value in (D) (shown by the distance between the tail and the head of the Waterfall plot) but very little information in (F).
ILIV-SHAP also tells the user whether an AI model’s prediction encodes useful complementary information from a particular feature. In Figure 3(E), the defendant’s prior count record (prior_counts=1) ranks second among features that impact the AI prediction ($-0.139$). Upon seeing this, users might be interested in investigating how the AI uses prior_counts=1. However, Figure 3(F) shows that prior_counts=1 contributes marginally to the decision-relevant information in AI predictions (about $+0.007$). Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.
@@ 7 Discussion and Limitations
We propose a decision-theoretic framework for assigning value to information in human-AI decision-making. Our methods quantify the additional information value of any signals over an agent’s decisions, and can be applied to any combination of artificial or human agent judgments. The three demonstrations we provide show how quantified information value can drive model selection, empirical evaluation, and explanation design. These are just a few of many possible use cases. For example, explanation approaches could be compared in terms of their information value to human decisions. Information value analysis could also drive elicitation of human private signals upon identifying that human decisions contain AI-complementary information. New explanation strategies based on ILIV-SHAP could more directly integrate visualized information value with conventional depictions of feature importance.
It is important to note that when using ACIV to understand complementary information over human decisions, it cannot be interpreted as definitively establishing that particular signals were used by the human. It is always possible that the human has other private signals offering equivalent information to a feature being analyzed. This raises a question of whether it is definitive about use if it is an AI as the agent instead of a human. However, quantifying the value of observed information is an important first step toward learning about information that may exist beyond an explicit problem definition.
Our framework quantifies the best-attainable performance improvement from integrating signals in decisions. This does not necessarily mean that highlighting or otherwise emphasizing those signals for humans will lead to them performing better. However, the basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021). For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the rational decision-maker to only make monotonic decisions with AI prediction and their own predictions.
@@ Appendix A The Combinatorial Nature of the Value of Signals
$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(\operatorname{ACIV}(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (10)
@@
The following algorithm provides a polynomial-time approximation of the Shapley value of $\operatorname{ACIV}$. Under the assumption of submodularity, it orders the signals the same as the Shapley value.
Algorithm 1 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
@@
4:     $\phi^{\prime}_{j}=\operatorname{ACIV}(\Sigma_{j};V^{*})\ for\ each\ j$
@@
Appendix B Embeddings for High-dimensional Signals in Decision Problem
Suppose the data-generating process is estimated from a set of observations $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
$\hat{\pi}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
In this case, the sample size $T$ that is required to get a good estimate of DGP is the exponential of the number of dimensions of the signal $v$. The high-dimensional signals, such as images and text, that are usual in human-AI decision problems are not computationally feasible to be quantified under our framework. We propose a systematic procedure to pick the optimal reduction model to represent the high-dimensional signal space in a low-dimensional embedding space.
The optimal reduction model and its generated embeddings should fulfill two properties. First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals. Second, the derived DGP that is implied by the embedding signals should be feasible to be estimated from the observed dataset. We quantify the above objectives by a framework based on Train/Test split.
Algorithm 2 Embedding high-dimensional signals in decision problems
1:  Input: Observed dataset $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Candidate model generator $C$.
2:  $R^{*}=0$
3:  Split the dataset into a training set and a testing set.
4:  repeat
5:     Generate a candidate model $c$ from $C$.
6:     Generate embedding $v^{c}_{i}=c(v_{i})$ on each data point of both sets.
7:     Estimate the DGP $\hat{\pi}$ implied by the new signals $v^{c}_{i}$ by Equation 11.
8:     Calculate rational decision rule:   $d^{r}(v^{c})=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v^{c})}\![{S(d,\omega)}]$.
9:     Calculate training performance $R_{training}=\text{\bf E}_{(v^{c},\omega)\sim training\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
10:     Calculate testing performance $R_{testing}=\text{\bf E}_{(v^{c},\omega)\sim testing\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
11:     if $R_{training}-R_{testing}>\tau$ then
12:        go to step 4.
13:     end if
14:     if $R^{*}<R_{training}$ then
15:        $c^{*}=c$.
16:        $R^{*}=R_{training}$.
17:     end if
18:  until no more new candidate model in $C$ or $R^{*}$ converages
19:  Output: $c^{*}$
The threshold $\tau$ represents the tolerance in the exchangeability between the signal-state tuples in the training set and the testing set. When the candidate model space is infinite, the optimal reduction model can still be found if the candidate space is a lattice that is defined by the informativeness of their generated embeddings, e.g., a class of embedding models with different parameter p in the dropout layer.
Appendix C Robustness Analysis in Experiment I
Figure 4: Robust analysis for experiment I on all V-shaped payoff functions. The kink $\mu$ is shown on the x-axis. Each subplot displays the difference between the $\operatorname{ACIV}$ on the row model over human decisions and the $\operatorname{ACIV}$ on the column model over human decisions. A positive value (colored in blue) at $\mu$ indicates the model on the row contains more informative than the model on the column under the evaluation of V-shaped scoring rule with kink $\mu$. The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
Appendix D More examples of ILIV-SHAP